# Novel Diffusion Models for Multimodal 3D Hand Trajectory Prediction

### v1

@@ III-A System Overview
Here we first provide the overall inference pipeline of our MMTwin in Fig. II. MMTwin receives multimodal data including egocentric RGB images =$\mathcal{I}=\{I_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(I_{t}\in\mathbb{R}^{c\times h\times w})$, point clouds $\mathcal{D}=\{D_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(D_{t}\in\mathbb{R}^{n\times 3})$, past 3D hand waypoints $\mathcal{H}_{p}=\{H_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(H_{t}\in\mathbb{R}^{3})$, and text prompt $\mathcal{O}$, and predicts future 3D hand trajectories $\mathcal{H}_{\text{f}}=\{H_{t}\}_{t=1}^{N_{\text{f}}}$. $N_{\text{p}}$ and $N_{\text{f}}$ denote the numbers of frames in the past and future time horizons respectively. $\mathcal{I}$ and $\mathcal{D}$ are both captured by headset RGBD camera. $\mathcal{O}$ is set as hand as proposed in the previous work [10]. Following [8, 7], we predict future hand waypoints in the global coordinate system, which is assigned as the camera coordinate system in the first frame of the input sequence $(t=-N_{\text{p}}+1)$.
MMTwin first calculates sequential homography =$\mathcal{M}=\{M_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(M_{t}\in\mathbb{R}^{3\times 3})$ from $\mathcal{I}$ as past camera egomotion following [9], and encodes them to past egomotion latents for the egomotion diffusion. $M_{t}$ represents the homography matrix between $t\,\text{th}$ frame and the first frame $(t=-N_{\text{p}}+1)$ estimated by SIFT descriptors [22] with RANSAC [23]. The egomotion diffusion predicts future egomotion latents conditioned on past ones, which will be further used as conditions for the HTP diffusion. The input images $\mathcal{I}$ and the text prompt $\mathcal{O}$ are fed into a foundation model [24] to generate visual semantic features. A fusion module proposed by the previous work [10] incorporates past hand waypoints and visual semantic features to generate HTP latents for the HTP diffusion. The input sequential point clouds $\mathcal{D}$ are first transformed to the unified global coordinate system by visual odometry, and then we voxelize the aggregated points into discrete voxel grids to reduce memory consumption and improve running efficiency. A voxel encoder is built based on 3D convolutions to convert the dense grids to 3D voxel patches. We exploit the vanilla Mamba [25] (VM) as the denoising model in the egomotion diffusion, while we propose a hybrid Mamba-Transformer module (HMTM) for denoising in the HTP diffusion. The denoised HTP latents are ultimately decoded to future 3D hand waypoints $\mathcal{H}_{\text{f}}$. As can be seen, MMTwin achieves decoupling predictions of camera egomotion and 3D hand trajectories, and bridges them through denoising conditions, following the fact that there is a synergy between hand movements and camera egomotion within the future interaction process.
@@ III-B Multimodal Feature Extraction
Vision encoder. As shown in Fig. II(a), we extract visual semantic features from past RGB images with text prompt hand following the previous work [10]. We use pretrained GLIP [24] here as the vision encoder. The visual grounding ability of GLIP enables our visual encoder to semi-implicitly capture hand poses and hand-scenario relationships within each 2D image frame, since the text prompt is used to indicate which part of the image should be focused on. Specifically, we extract the outputs of the deepest cross-modality multi-head attention module (X-MHA) in GLIP, which are denoted as $X^{\text{sem}}\in\mathbb{R}^{(N_{\text{p}}+L)\times x}$. $x$ is the feature channel dimension. $L$ equals $N_{\text{f}}$ during training, and is set to 00 during inference since future HTP latents will be replaced by sampled noises (noisy future latents in Fig. II(a)) in the inference process of our HTP diffusion.
@@ IV-A Experimental Setups
MMTwin configuration. We voxelize input point clouds into $20\times 20\times 20$ grids with the resolution of $0.05$ m. The voxel patches are with the size of $27\times 1024$. For both twin diffusion models, we set the channel dimension of the latent features to $1024$. The total number of diffusion steps is set to $1000$, while the egomotion diffusion takes only one step to predict egomotion features for high efficiency and the HTP diffusion takes 100 steps to predict future HTP features. Egomotion-aware Mamba blocks of HMTM are with convolutional kernel size $d\_conv\!=\!2$, hidden state expansion $expand\!=\!1$, and hidden dimension $d\_state\!=\!16$. The number of heads in the structure-aware Transformer of HMTM is set to $n\_head=4$, and the intermediate dimension of the feed-forward layer is $d\_f\mkern-2.0muf\mkern-2.0mun\!=\!2048$. We train MMTwin using AdamW optimizer [32] with a learning rate of 5e-5 for 1K epochs on EgoPAT3D-DT, H2O-PT, and our self-recorded datasets, and with a learning rate of 5e-6 for 2K epochs on the HOT3D-Clips dataset. Training and inference are both operated on 2 NVIDIA A100 GPUs.

### v2

@@ III-A System Overview
Here we first provide the overall inference pipeline of our MMTwin in Fig. II. MMTwin receives multimodal data including egocentric RGB images $\mathcal{I}=\{I_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(I_{t}\in\mathbb{R}^{c\times h\times w})$, point clouds $\mathcal{D}=\{D_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(D_{t}\in\mathbb{R}^{n\times 3})$, past 3D hand waypoints $\mathcal{H}_{p}=\{H_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(H_{t}\in\mathbb{R}^{3})$, and text prompt $\mathcal{O}$, and predicts future 3D hand trajectories $\mathcal{H}_{\text{f}}=\{H_{t}\}_{t=1}^{N_{\text{f}}}$. $N_{\text{p}}$ and $N_{\text{f}}$ denote the numbers of frames in the past and future time horizons respectively. $\mathcal{I}$ and $\mathcal{D}$ are both captured by headset RGBD camera. $\mathcal{O}$ is set as hand as proposed in the previous work [10]. Following [8, 7], we predict future hand waypoints in the global coordinate system, which is assigned as the camera coordinate system in the first frame of the input sequence $(t=-N_{\text{p}}+1)$.
MMTwin first calculates sequential homography $\mathcal{M}=\{M_{t}\}_{t=-N_{\text{p}}+1}^{0}\,(M_{t}\in\mathbb{R}^{3\times 3})$ from $\mathcal{I}$ as past camera egomotion following [9], and encodes them to past egomotion latents for the egomotion diffusion. $M_{t}$ represents the homography matrix between $t\,\text{th}$ frame and the first frame $(t=-N_{\text{p}}+1)$ estimated by SIFT descriptors [22] with RANSAC [23]. The egomotion diffusion predicts future egomotion latents conditioned on past ones, which will be further used as conditions for the HTP diffusion. The input images $\mathcal{I}$ and the text prompt $\mathcal{O}$ are fed into a foundation model [24] to generate visual semantic features. A fusion module proposed by the previous work [10] incorporates past hand waypoints and visual semantic features to generate HTP latents for the HTP diffusion. The input sequential point clouds $\mathcal{D}$ are first transformed to the unified global coordinate system by visual odometry, and then we voxelize the aggregated points into discrete voxel grids to reduce memory consumption and improve running efficiency. A voxel encoder is built based on 3D convolutions to convert the dense grids to 3D voxel patches. We exploit the vanilla Mamba [25] (VM) as the denoising model in the egomotion diffusion, while we propose a hybrid Mamba-Transformer module (HMTM) for denoising in the HTP diffusion. The denoised HTP latents are ultimately decoded to future 3D hand waypoints $\mathcal{H}_{\text{f}}$. As can be seen, MMTwin achieves decoupling predictions of camera egomotion and 3D hand trajectories, and bridges them through denoising conditions, following the fact that there is a synergy between hand movements and camera egomotion within the future interaction process.
@@ III-B Multimodal Feature Extraction
Vision encoder. As shown in Fig. II(a), we extract visual semantic features from past RGB images with text prompt hand following the previous work [10]. We use pretrained GLIP [24] here as the vision encoder. The visual grounding ability of GLIP enables our visual encoder to semi-implicitly capture hand poses and hand-scenario relationships within each 2D image frame, since the text prompt is used to indicate which part of the image should be focused on. Specifically, we extract the outputs of the deepest cross-modality multi-head attention module (X-MHA) in GLIP, which are denoted as $X^{\text{sem}}\in\mathbb{R}^{(N_{\text{p}}+L)\times x}$. $x$ is the feature channel dimension. $L$ equals $N_{\text{f}}$ during training, and is set to 0 during inference since future HTP latents will be replaced by sampled noises (noisy future latents in Fig. II(a)) in the inference process of our HTP diffusion.
@@ IV-A Experimental Setups
MMTwin configuration. We voxelize input point clouds into $20\times 20\times 20$ grids with the resolution of $0.05$ m. The voxel patches are with the size of $27\times 1024$. For both twin diffusion models, we set the channel dimension of the latent features to $1024$. The total number of diffusion steps is set to $1000$, while the egomotion diffusion takes only one step to predict egomotion features for high efficiency and the HTP diffusion takes 100 steps to predict future HTP features. Egomotion-aware Mamba blocks of HMTM are with convolutional kernel size d_conv=2d\_conv\!=\!2, hidden state expansion expand=1expand\!=\!1, and hidden dimension d_state=16d\_state\!=\!16. The number of heads in the structure-aware Transformer of HMTM is set to n_head=4n\_head=4, and the intermediate dimension of the feed-forward layer is $d\_f\mkern-2.0muf\mkern-2.0mun\!=\!2048$. We train MMTwin using AdamW optimizer [32] with a learning rate of 5e-5 for 1K epochs on EgoPAT3D-DT, H2O-PT, and our self-recorded datasets, and with a learning rate of 5e-6 for 2K epochs on the HOT3D-Clips dataset. Training and inference are both operated on 2 NVIDIA A100 GPUs.